# Revisiting kink-like parametrization and constraints using OHD/Pantheon/BAO samples

### v1

@@ IV.3 When $\alpha=1/3$
We also considered the case of $\Lambda$CDM, when $\alpha$ is also fixed, i.e. $\alpha=1/3$. The parameter space here reduces to $(H_{0},z_{t})$. We have obtained $q_{0}$ from the same equation (6), mentioned in Table 1. We present the $68\%$ and $95\%$ confidence levels for this parameter space in FIG. 3. It is clear that the $\Lambda$CDM is in good agreement with the data and used for comparing with the two previous cases. Here, the BAO measurements are giving us tighter constraints on the parameters. The behavior of $q(z)$ shows a signature flip and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase in the recent past and approaches a de-sitter phase as expected. In this case, the jerk parameter is $j=1$ for the best-fit values, consistent with the observations.
Figure 3: The 1-dimensional and 2-dimensional marginalized posterior distributions in the parameter space $(H_{0},z_{t})$ for the case when $\alpha=1/3$.
In FIG. 4, we have shown the evolution of the Hubble parameter $H(z)$ as a function of redshift $z$ compared with 31 $H(z)$ measurements. We have plotted the data points for $H(z)$ with $1\sigma$ error bars. We have observed that our model is well consistent with the $H(z)$ data against redshift for free $q_{f}$, $q_{f}=-1$, $\alpha=1/3$. For a convenient display, we also show $H(z)/(1+z)$ in FIG. 4.
Figure 4: Figure shows the evolution of Hubble parameter $H(z)$ and $H(z)/(1+z)$ as a function of $z$. We have plotted 31 $H(z)$ measurements with its $1\sigma$ error bars for all the cases considered.
@@
OHD $68.14\pm 0.11$ $-0.522^{+0.037}_{-0.034}$ $0.623\pm 0.060$ $1/3$ $18.193$ $22.193$ $25.061$
$\alpha$ fixed SN $68.13\pm 0.10$ $-0.545^{+0.027}_{-0.028}$ $0.663\pm 0.049$ $1/3$ $1044.981$ $1048.981$ $1058.89$
OHD+SN+BAO $68.129\pm 0.099$ $-0.54^{+0.022}_{-0.023}$ $0.654\pm 0.039$ $1/3$ $1066.039$ $1078.039$ $1090.4$
Table 1: Summary of the best-fit values of model parameters and statistical analysis using OHD, SN, and OHD+SN+BAO, including the 68% and 95% confidence levels.
Figure 5: The plots represents the present values of $H_{0}$ with their corresponding error bars obtained from several works. The green, blue, and black lines represent the best-fit estimated values obtained for the current model in all three scenarios (free $q_{f}$, $q_{f}=-1$, $\alpha=1/3$), with the related $1\sigma$ error shown as shaded portions [40].
@@ V Conclusion
In this work, we reexamined a kink-like parametrization for the deceleration parameter to investigate the transition from cosmic deceleration to acceleration in a spatially flat FLRW universe. The functional form of $q(z)$ involves $q_{f}$, $q_{i}$, $q_{0}$, $z_{t}$, $\alpha$, and $H_{0}$ as free parameters. Given that our parametrization specifically intends to explain the cosmic evolution that begins with a matter-dominated decelerated phase, we set the value of $q_{i}=1/2$. However, as new data pour in and new techniques evolve, revisiting the nature of $q$ with newer datasets is imperative. So, here, we focussed on recent OHD, SN (Pantheon samples), and BAO/CMB observations and carried out a Markov chain Monte Carlo (MCMC) analysis of the model, which allowed us to determine the optimal values of the model parameters.
We first consider the case in which we do not impose a fixed value for $q_{f}$. So, we constrained four parameters namely $(H_{0},q_{0},z_{t},\alpha)$. The comparison of the observational data on the OHD shows a good agreement with the $\Lambda$CDM model. The AIC analysis also substantiates the presence of a little tension for $OHD+SN+BAO$ between the current model for this scenario and the $\Lambda$CDM predictions. However, to definitively address this matter, additional observational evidence covering a more comprehensive range of redshifts is required. There is an uninspired phase of the universe when $z\rightarrow-1$ since $q(z)$ diverges more from the de-sitter phase for the combined dataset. Another choice for $q_{f}$ could also be made to ensure the model’s consistency in the distant future. The evolution of the equation of state parameter reveals that the $\omega$ is positive at higher redshifts and remains greater than $-1$ (quintessence phase) until the present time but crosses a phantom divide for the far future. The present value of $\omega$ is $-0.72$, $-0.78$, and $-0.76$ for $OHD$, $SN$, and $OHD+SN+BAO$, respectively. The current value of $j$ exceeds $1$ for all the datasets ($j_{0}=1.3$, $1.71$, $1.7$ for $OHD$, $SN$, $OHD+SN+BAO$, respectively). But in this case, the $j(z)$ is not well constrained as it may be connected to the appearance of abrupt future singularities.
We also examined the scenario where the model exhibits the final de-Sitter phase, characterized by $q_{f}=-1$. The Pantheon samples are the primary source for improvements compared to prior works. It is anticipated that these samples lead to a substantial enhancement in the constraints, owing to their higher statistics. By incorporating these datasets, we analyze the influence of $H_{0}$ on the transition $z_{t}$. We find that the cosmological deceleration-acceleration transition depends on the value of $H_{0}$ but is otherwise mildly dependent on other cosmological parameters. The Hubble parameter is not so far from the current bounds, agreeing with the Scolnic 2018 results [23]. Here, the jerk parameter is well constrained as it converges to $j_{0}=1$ at late times.
@@
Our results have shown that the present model may represent an interesting alternative to dark energy for the case of $q_{f}=-1$ to tighten the constraints on the parameters for $OHD$, $Pantheon$, and $BAO$. Furthermore, it is worth noting the potential of future surveys such as DESI, LSST, SH0ES, and BINGO observations in effectively constraining the range of these model parameters.
@@ Acknowledgments
SA & PKS thank IUCAA, Pune, India, for the hospitality where part of the work was carried out. PKS acknowledges the Science and Engineering Research Board, Department of Science and Technology, Government of India, for financial support to carry out Research Project No.: CRG/2022/001847.

### v3

@@ IV.3 When $\alpha=1/3$
We also considered the case of $\Lambda$CDM, when $\alpha$ is also fixed, i.e. $\alpha=1/3$. The parameter space here reduces to $(H_{0},z_{t})$. We have obtained $q_{0}$ from the same equation (6), mentioned in Table 1. We present the $68\%$ and $95\%$ confidence levels for this parameter space in FIG. 5. It is clear that the $\Lambda$CDM is in good agreement with the data and used for comparing with the two previous cases. Here, the measurements are giving us tighter constraints on the parameters. The behavior of $q(z)$ shows a signature flip and $z_{t}$ indicates the epoch when the expansion of the universe went from a decelerating to an accelerating phase in the recent past and approaches a de-sitter phase as expected. In this case, the jerk parameter is $j=1$ for the best-fit values, consistent with the observations (see FIG. 6).
In FIG. 7, we have shown the evolution of the Hubble parameter $H(z)$ as a function of redshift $z$ compared with 31 $H(z)$ measurements. We have plotted the data points for $H(z)$ with $1\sigma$ error bars. It is observed that our model is well consistent with the $H(z)$ data against redshift for free $q_{f}$, $q_{f}=-1$, $\alpha=1/3$. Furthermore, the behavior of $\mu(z)$ (for Pantheon+) is seen in FIG. 7 for all three cases.
Figure 5: The marginalized posterior distributions in the parameter space $(H_{0},z_{t})$ for the case when $\alpha=1/3$.
Figure 6: Behavior of the jerk parameter when $\alpha=1/3$ for $SN$ and $OHD+SN+BAO$.
@@
$\alpha$ fixed SN $72.459\pm 0.092$ $-0.441^{+0.022}_{-0.021}$ $0.500\pm 0.031$ $1/3$ $1609.89$ $1613.89$ $1624.76$
Joint $72.530\pm 0.086$ $-0.486\pm 0.017$ $0.566\pm 0.027$ $1/3$ $1642.834$ $1646.834$ $1657.75$
Table 1: Summary of the best-fit values of model parameters and statistical analysis using SN and OHD+SN+BAO, including the confidence levels.
Figure 7: The evolution of Hubble parameter $H(z)$ and $\mu(z)$ as a function of $z$. We have plotted $H(z)$ measurements and Pantheon+ measurements with its error bars for all the cases.
Figure 8: The plots represent the present values of $q_{0}$ and $H_{0}$ with their corresponding error bars obtained from several works. The green, blue, and black lines represent the best-fit estimated values obtained for the current model in all three scenarios (free $q_{f}$, $q_{f}=-1$, $\alpha=1/3$), with the related $1\sigma$ error shown as shaded portions [36, 37].
@@ V Conclusion
In this work, we reexamined a kink-like parametrization for the deceleration parameter to investigate the transition from cosmic deceleration to acceleration in a spatially flat FLRW universe. The functional form of $q(z)$ involves $q_{f}$, $q_{i}$, $q_{0}$, $z_{t}$, $\alpha$, and $H_{0}$ as free parameters. Given that our parametrization specifically intends to explain the cosmic evolution that begins with a matter-dominated decelerated phase, we set the value of $q_{i}=1/2$. However, as new data pour in and new techniques evolve, revisiting the nature of $q$ with newer datasets is imperative. So, here, we focussed on recent OHD, SN (Pantheon+ samples), and BAO observations and carried out an MCMC analysis of the model, which allowed us to determine the optimal values of the model parameters.
We first consider the case in which we do not impose a fixed value for $q_{f}$. So, we constrained four parameters namely $(H_{0},q_{0},z_{t},\alpha)$. The comparison of the observational data on the $SN$ shows a good agreement with the $\Lambda$CDM model. The AIC analysis also substantiates the presence of a little favor for $OHD+SN+BAO$ between the current model for this scenario and the $\Lambda$CDM predictions. However, to definitively address this matter, additional observational evidence covering a more comprehensive range of redshifts is required. There is an uninspired phase of the universe when $z\rightarrow-1$ since $q(z)$ diverges more from the de-sitter phase for these datasets. Another choice for $q_{f}$ could also be made to ensure the model’s consistency in the distant future. The evolution of the equation of state parameter reveals that the $\omega$ is positive at higher redshifts and remains greater than $-1$ (quintessence phase) until the present time but crosses a phantom divide for the far future. The present value of $\omega$ is $-0.532^{+0.021}_{-0.018}$ and $-0.54\pm 0.13$, for $SN$ and $OHD+SN+BAO$, respectively. Observations indicate that, in this model, the current value of $j$ is far from $1$ for all the datasets ($j_{0}=0.05$ and $0.02$ for $SN$, $OHD+SN+BAO$, respectively). But in this case, the $j(z)$ is not well constrained as it may be connected to the appearance of abrupt future singularities.
We also examined the scenario where the model exhibits the final de-Sitter phase, characterized by $q_{f}=-1$. The Pantheon+ samples are the primary source for improvements compared to prior works. It is anticipated that these samples lead to a substantial enhancement in the constraints, owing to their higher statistics. By incorporating these datasets, we analyze the influence of $H_{0}$ on the transition $z_{t}$. We find that the cosmological deceleration-acceleration transition depends on the value of $H_{0}$ and is also mildly dependent on other cosmological parameters. Here, the jerk parameter is well constrained as it converges to $j_{0}=1$ at late times.
@@
Our results have shown that the present model may represent an interesting alternative to dark energy for the case of $q_{f}=-1$ to tighten the constraints on the parameters for $OHD$, $Pantheon+$, and $BAO$. Our current research reveals a deviation of our model from the standard $\Lambda$CDM when considering a combined dataset, indicating a preference for utilizing a parametrization of the dynamical parameter. As a result, there is an opportunity to delve deeper into understanding cosmological tensions and their resolutions, necessitating the development of new cosmological probes beyond our current study. In the future, we aim to incorporate new observational datasets such as BINGO, LSST, and DESI to investigate cosmological tensions further.
@@ Acknowledgments
SA & PKS thank IUCAA, Pune, India, for the hospitality where part of the work was carried out. PKS acknowledges the Science and Engineering Research Board, Department of Science and Technology, Government of India, for financial support to carry out Research Project No.: CRG/2022/001847. We are very much grateful to the honorable referee and to the editor for the illuminating suggestions that have significantly improved our work in terms of research quality and presentation.